A horizontal axis wind turbine has a rotor diameter of 76 m and reaches its full load output at the generator terminals of 2 MW at a wind speed of 12 m/s. The overall efficiency of the gearbox and generator is 94 %. Assume the density of air is 1.2 kg/m3 and calculate the power coefficient (Cp) of the aerodynamic rotor [6]


1
Expert's answer
2021-10-10T09:32:03-0400

Power produced by wind turbine,P="\frac{1}{2}\rho Av^{3}C_p\eta"

P=2*106 W

v=12m/s

"\eta" =0.94

"\rho" =1.2kg/m3

r="\frac{76}{2}=38m"

"2*10^{6}=\frac{1}{2}(12)(π*38^{2})*12^{3}C_p*0.94"

"C_p=0.452"